1. Mayan Company had net income of $34,830. The weighted-average common shares outstanding were 8,600. The company has no preferred stock. The company's earnings per share is:

Multiple Choice

  • $4.11.

  • $3.99.

  • $5.00.

  • $1.17.

  • $4.05.

2. A company has net income of $875,000; its weighted-average common shares outstanding are 175,000. Its dividend per share is $1.20, its market price per share is $103, and its book value per share is $98.50. Its price-earnings ratio equals:

Multiple Choice

  • 4.50.

  • 20.60.

  • 19.70.

  • 3.30.

  • 5.70.

3. A company paid $0.56 in cash dividends per share. Its earnings per share is $4.28 and its market price per share is $29.00. Its dividend yield equals:

Multiple Choice

  • 14.76%.

  • 1.93%.

  • 13.08%.

  • 7.64%.

  • 5.18%.
Answer #1
1.
Earnings per share = Net income / Weighted average common shares outstanding = 34830 / 8600 4.05
2.
Earnings per share = Net income / Weighted average common shares outstanding = 875000 / 175000 5
Price-earnings ratio = Market price per share / Earnings per share = 103 / 5 20.60
3.
Dividend yield = Cash dividend per share / Market price per share = 0.56 / 29.00 1.93%